OpenStudy (anonymous):

Algebra Help: Find the roots of each equation by factoring 12x=9x^2+4 16x^2=9

OpenStudy (anonymous):

To factor an expression, you first want all the terms on one side of the '=' sign: \[ 0 = 9x^2-12x+4 \] Now, you're looking for a factoring that looks roughly like \[ (ax\pm b)(cx \pm d)\] So, what numbers multiply together to get 9 and what other numbers multiply together to get 4?

OpenStudy (anonymous):

3 and 3

OpenStudy (anonymous):

2 and 2

OpenStudy (anonymous):

Great! One more thing, when looking at expressions like \[ 9x^2 -12x+4\] the second operation tells us whether the pluses and minuses in \[(ax\pm b)(cx \pm d)\] will be the same or different. We have a '+', so we know the signs are going to be the same. And then '-' before 12x tells us that both those signs are going to be minuses. Test it out! Does \[ (3x-2)(3x-2) = 9x^2 -12 + 4\] ?

OpenStudy (anonymous):

Yeah it does.

OpenStudy (anonymous):

Would that be the root?

OpenStudy (anonymous):

Cool. The roots of the equation are the values of x that make the function 0. So, we've turned \[ 9x^2-12x+4 = 0 \] into \[(3x-2)(3x-2) = 0 \] For what value of x does the above equation hold true?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Not quite. When x = 1, \[ (3x-2)(3x-2) = (3-2)(3-2) = 1 \] Think of it this way: you're multiplying two numbers and getting a 0. That means one of those numbers has to be equal to zero, right? So, how do you solve 3x-2= 0?

OpenStudy (anonymous):

.66

OpenStudy (anonymous):

thats what I got.

OpenStudy (anonymous):

Correct! So 2/3 is a root of 12x= 9x^2-4

OpenStudy (anonymous):

Is there 2 roots?

OpenStudy (anonymous):

Or just the one?

OpenStudy (anonymous):

There are two roots, but in this case both roots happen to have the value 2/3.

OpenStudy (anonymous):

If you had two different numbers. Would you do it with both of the equations?

OpenStudy (anonymous):

Also. How do I do the second one?

OpenStudy (anonymous):

I'm not sure I understand your first question. If you had two different roots (say a and b), your equation would look something like \[ (x-a)(x-b) = 0 \] Try repeating the same process for the second one. Put all the x's and numbers on one side of the equal sign. And then think of multiples

OpenStudy (anonymous):

Okay. Thats what I meant. Can you help me factor the second one?

OpenStudy (anonymous):

Glad I understood your question then haha Sure! So first, rewrite the equation into \[ 16x^2 - 9 = 0 \] Again, you're looking for ways to break up 16 and 9

OpenStudy (anonymous):

4 and 4 for sixteen and 3 and 3 for nine So you find the x for both roots?

OpenStudy (anonymous):

both equations

OpenStudy (anonymous):

Which equations are you talking about?

OpenStudy (anonymous):

so if you factor 16 and 9, you get (x+4)^2=0 and (x+3)^3=0 right?

OpenStudy (anonymous):

Not quite, beacause \[ (x+4)(x+3) \neq 16x^2-9\] Remember that the factors you find for 16 go into the a/c positions and the factors of 9 go into the b/d positions in: \[ (ax\pm b)(cx \pm d)\]

OpenStudy (anonymous):

oh. Okay so it would be -3/4

OpenStudy (anonymous):

That's one of the roots! Now the second one...

OpenStudy (anonymous):

Second one?

OpenStudy (anonymous):

The equation has 2 roots which are different. When you factored, you should have gotten \[ (4x-3)(4x+3) \]

OpenStudy (anonymous):

the other would be 3/4

OpenStudy (anonymous):

Can you help with another thing?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

Write a quadratic function in standard form for each given set of zeros. -2 and 7 1 and -8

OpenStudy (anonymous):

Ah, so it's the same process, but in reverse. You're given the roots and need the equation. For -2, you're looking for some expression involving an x so that when x = -2, the expression will equal 0. Does that make sense?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

Haha okay. Let me use your first question as an example. When we were looking for roots, we had \[ 3x-2 = 0 \]right? And we used that to solve for x and get x=2/3. From this equation, you can also work backwards to the first! x=2/3 3x = 2 (3x - 2) = 0 Does that make sense?

OpenStudy (anonymous):

Yeah. But I dont know how to get the other two numbers.

OpenStudy (anonymous):

What 2 numbers are you talking about? Now you're given -2 and 7. Let's look at the 7. This means there's a root at x=7. x=7 (x-7) = 0 If you do the same for x=-2, you get (x+2) = 0. So your expression looks like (x+2)(x-7) = 0

OpenStudy (anonymous):

Oh. WHat do I do next? Or is that it?

OpenStudy (anonymous):

Well, the problem statement says "quadratic function in standard form" so you're probably supposed to multiply the two terms together and get something like x^2+_x+_ = 0

OpenStudy (anonymous):

(by probably, I mean you are xD)

OpenStudy (anonymous):

Can you help with the other one? I got x^2-5x-14

OpenStudy (anonymous):

For the one we just did

OpenStudy (anonymous):

You're correct about the first one! Why don't you take the lead on this question and I'll tell you how you're doing? It's very similar to the one we just did. You're starting off with x=1 and x=-8.

OpenStudy (anonymous):

(x+1)(x-8)=0

OpenStudy (anonymous):

Very close. You've got your signs switched for both! Remember that x=-8 ---> x+8 = 0

OpenStudy (anonymous):

Oh. Then you would get x^2+7x-8

OpenStudy (anonymous):

Yay! Yes, you're right.

OpenStudy (anonymous):

Okkay. Thank you. That's all I needed help with.

OpenStudy (anonymous):

You're very welcome. I'm happy to help :)